Circular No. 7865

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002by IN ESO 139-G34 R. Chassagne, Ste. Clotilde, Île de Réunion, reports his discovery of

an apparent supernova (mag 14.5) on an unfiltered CCD frame taken on Mar. 31.02 UT (and confirmed on frames taken on Apr. 2.04) with a 0.30m automated telescope. SN 2002by is located at α = 17^{h}50^{m}18^{s}.58, δ = −59^{o}33′51′′.6 (equinox 2000.0), which is 24′′.9 west and 13′′.4 south of the center of ESO 139-G34. The new object does not appear on unfiltered images taken by Chassagne on Feb. 25 (limiting mag ∼ 18) or on a U.K. Schmidt plate taken on 1975 June 14 (limiting mag 21.0).

POSSIBLE SUPERNOVA IN NGC 7714 S. Mattila and P. Meikle, Imperial College, London; N. Walton, In-

stitute of Astronomy, Cambridge; R. Greimel, Isaac Newton Group, La Palma; S. Ryder, Anglo-Australian Observatory; and C. Alard, DASGAL and Institut d’Astrophysique, Paris, report the discovery (7σ detection) of a possible supernova (mag 17.3) in an archival U.K. Infrared Telescope (UKIRT) K-band image obtained with IRCAM3 on 1998 Sept. 5.4 UT; images taken 3 min apart showed no detectable movement (seeing 0′′.8). The new object was located in the nuclear region of the starburst galaxy NGC 7714 at α = 23^{h}36^{m}14^{s}.0, δ = +2^{o}09′24′′ (equinox 2000.0), which is 2′′ west and 5′′ north of (or 1 kpc from) the galaxy nucleus. The new object was apparent via image-subtraction techniques when compared with later K_{s}- band images obtained on 2001 Sept. 1.0 at the William Herschel Telescope (+ INGRID) and on 2000 Oct. 21 by A. Lancon and collaborators at the New Technology Telescope (+ SOFI), when no point source was visible at the location. Subtraction techniques reveal no object (limiting mag 18.5) at this location on a UKIRT H-band image obtained on 1998 Sept. 5 (NTT reference image). If this was a normal core-collapse supernova at any epoch, or a slow decliner (Mattila and Meikle 2001, MNRAS 324, 325) at an early epoch, an estimate of the H − K color limit implies an A_{V} towards the supernova of at least 6 mag. Alternatively, such a red H − K color could be produced by a slow-decliner type (e.g., SN 1998S) at a late phase with a much lower A_{V}. The search of archives for other detections of this possible supernova is encouraged.

W. Li, University of California, finds no object at the location of the possible supernova on KAIT images from 1998 Oct. 22 (limiting mag 19.0), and T. Puckett, Mountain Town, GA, finds nothing on his image taken on 2000 Nov. 11.16 that is not present on a red Digitized Sky Survey image.

2002 April 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green